OpenStudy (anonymous):

Solve for x. 4 • 7 – 4 • 10 = 4(x – 10)

OpenStudy (anonymous):

x = 70

OpenStudy (anonymous):

That was easy. :P

OpenStudy (anonymous):

Medal?

OpenStudy (anonymous):

WRONG

OpenStudy (anonymous):

Click over there ---->

OpenStudy (anonymous):

These are the answer choice: A.-13 B.-6 C. 4 D. 7

OpenStudy (anonymous):

WTF

OpenStudy (anonymous):

The answer is -6. Medal?

OpenStudy (anonymous):

Incorrect. Test answer = 7

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

\[28-40=4(x-10),\frac{ -12 }{ 4 }=x-10, calculate ~x\]

OpenStudy (anonymous):

28 - 40 = 4 x - 40 28 = 4 x x = 7

OpenStudy (anonymous):

x=7

OpenStudy (j2lie):

the answer is 7 because 4*7-4*10=28-40=-12. When you minus negative ten and positive you get four and when you multiply it by -2 you get -12.

OpenStudy (anonymous):

x=7